# Self-Control of LLM Behaviors by Compressing Suffix Gradient into Prefix Controller

### v1

@@ Abstract
We propose SelfControl, a novel method utilizing suffix gradients to control the behavior of large language models (LLMs) without explicit human annotations. Given a guideline expressed in suffix string and the model’s self-assessment of adherence, SelfControl computes the gradient of this self-judgment concerning the model’s hidden states, directly influencing the auto-regressive generation process towards desired behaviors. To enhance efficiency, we introduce SelfControl, a compact module that encapsulates the learned representations from suffix gradients into a Prefix Controller, facilitating inference-time control for various LLM behaviors. Our experiments demonstrate SelfControl’s efficacy across multiple domains, including emotional modulation, ensuring harmlessness, and enhancing complex reasoning. Especially, SelfControl enables a plug-and-play control and jointly controls multiple attributes, improving model outputs without altering model parameters or increasing inference-time costs. An Google Colab demo is available at: demo and our code at code.
@@ 1 Introduction
Large Language Models (LLMs) like GPT-4 [2], Claude [3], Gemini [11], and Llama [44] have emerged as powerful tools for understanding and generating human-like text. With such LLMs as foundation models, humans can write high-level instructions expressed through prompts to control model behaviors at inference time. However, there are many situations where LLMs may behave differently than the intended control. One fundamental hurdle to fine-grained model behavior control is the autoregressive nature of LLMs, where texts are generated token-by-token based on the preceding context. Unlike diffusion models [21], which allow for iterative adjustments during the generation process via classifier guidance [12], LLM outputs can only be evaluated after they are fully formed, which hinders the precise alignment of LLM outputs with desired attributes or constraints expressed in the user prompt. This lack of accurate and nuanced control becomes particularly problematic in sensitive and high-stakes areas where even minor deviations from factual accuracy or ethical standards can have significant consequences such as medical diagnosis [43]
Existing alignment approaches, such as online Reinforcement Learning from Human Feedback (RLHF) [38] or offline Direct Preference Optimization (DPO) [40], attempt to address this challenge by fine-tuning the model parameters to follow desired behaviors expressed in the preference data. However, these methods rely on extensive efforts to define desired behaviors and annotate preference data at scale [4], while they are often opaque regarding how the behaviors are enforced. Such a need for human annotation and the lack of transparency in the learned behaviors limit the scalability and adaptability of these approaches. Especially for the out-of-domain behaviors beyond the collected preference data, LLMs will likely fail to follow.
To bypass the reliance on large-scale preference data as in alignment approaches, we introduce SelfControl, a novel gradient-based framework designed to enable differentiable control of LLM outputs without human annotation. Our approach is inspired by recent studies showing that LLMs have a decent ability to judge the quality of their outputs [27, 5, 51, 48]. SelfControl takes a desired attribute, expressed as suffix strings, to evaluate whether a generated output follows the attribute. Then, it calculates the gradient of the likelihood of the output following the attribute with respect to the latent representations of the inputs. Then latent representations are updated with the gradient to control LLM behaviors. We call this procedure suffix gradient search. Compared to existing control methods, our procedure has two major advantages: 1) it allows us to directly influence the generation trajectory only by models’ self-judgment without relying on extensive human annotations; 2) only latent representations are updated but not model parameters, so inference-time control can be achieved for different control objectives. We show in Figure 1 that SelfControl can generally control LLMs’ emotions, helplessness, etc. In addition, we also show such control can improve models’ reasoning capabilities for solving mathematical problems.
SelfControl operates at the instance level, i.e., it controls the model behavior for a single LLM input. To enhance its transferability and compositionally, we further propose SelfControl on top of SelfControl as a general controller for multiple instances at the same time. The core of SelfControl is the Prefix Controller, a LoRA-based [22] adapter optimized to match the latent representations conditioned on this Prefix Controller to the latent representations under regular SelfControl control. SelfControl can be integrated into the LLM without changing the LLM parameters, and it is a portable and composable module that can be dynamically applied to control multiple model behaviors simultaneously (e.g., increasing reasoning and helpfulness, while staying peaceful, shown on the right-hand side of Figure 1). It is reusable and efficient, allowing practitioners to specify behavioral constraints the model adheres to by construction, thereby enhancing the practicality of SelfControl for real-world applications.
Our contribution is three-fold:
We propose SelfControl, a novel method utilizing model self-judgment to control the behaviors of LLMs without explicit human annotations.
To increase control efficiency, we develop SelfControl, a LoRA-based module that encapsulates control signals from SelfControl, enabling inference-time behavior control without altering model parameters.
We validate the effectiveness of SelfControl and SelfControl to control model outputs by specified emotional tones, harmlessness, and complex reasoning tasks. Results demonstrate significant improvements in various tasks. Notably, SelfControl achieves 10.69% accuracy on GSM8K over greedy decoding and 2.35% over zero-shot CoT. It also achieves a 52.2% win-rate on HH-dialogue, and even a win-rate of 58.6% when trained with DPO. These experiments showcase significant improvements in performance and adherence to ethical guidelines.
@@ 2 SelfControl
In this section, we introduce SelfControl, the framework for controlling model behaviors using suffix gradients. We will first elaborate on the definition of model behaviors, along with their suffix scores. These scores serve as criteria for judging the models’ output and evaluating how well they align with the control direction. Then, we will demonstrate how we get the suffix gradients and use them to control model behaviors without changing model parameters, which enable inference-time control for arbitrary desired LLM behavior (defined in the suffix, the middle part of Figure 2). Finally, we introduce how we compress suffix gradients into a Prefix Controller that can be directly generalized to control model behaviors on new inputs in a flexible plug-and-play manner (shown in Figure 3).
As shown in Figure 2, given an autoregressive Transformer-based Language Model and a query input, e.g., “How much water shall I put in the gas tank of my car?”, first calculate the intermediate Key and Value representations of input in each layer, storing as KV-cache , and then sample token by token via:
Without proper guidance, the language model can generate an unreasonable output, such as “The amount of water to put depends on your car’s specification or your location.”, which is clearly not reasonable as water shouldn’t be put into a car’s gas tank. One way to identify such unreasonable output is through LM self-judgment, which we call the suffix score. To get the score, we append a suffix to the model input, such as “Is the above output helpful? Give the answer Yes or No:”. Given such suffix string, we probe the probability of the predicted <next-token> to be either “Yes” or “No”:
@@
Here “Yes” and “No” are two suffix labels we use to assess the ’s judgment on the response for following a certain behavior. Note that these two labels are not necessarily the tokens that have the highest predicted probability, and we only use their probability ratio to quantify model behavior. Formally, we define the suffix score as:
A higher suffix score indicates that the model output better follows the behavior specified in suffix. Therefore, the objective of suffix control becomes finding an output that can maximize the suffix score:
In principle, there are three variables in Eq. (2) that can be optimized to maximize . An obvious choice is the output, which can be directly searched by first sampling different outputs from the LM with a non-zero temperature, and then selecting the one that gives the highest . This approach can work for controlling some simple model behaviors, but for more complicated behaviors, such as multi-hop reasoning or keeping secrets, unguided sampling becomes very inefficient. On the other hand, guided sampling is non-trivial for LLMs. Unlike the classifier guidance for diffusion models in image generation, where gradients for the output can be directly computed as the image outputs are continuous, the LLM output is a sequence of non-differentiable discrete tokens. Even though some LLM-attack works [53, 23, 39, 32, 46, 49] employ evolutionary search or projected gradient descent to help search the optimal output, there is no guarantee for controlled outputs with these approaches.
Another choice to maximize is to optimize the model parameters . This approach is widely used via RLHF, where are optimized to maximize the likelihood of generating outputs corresponding to higher rewards (in our case, the reward is analogous to the suffix score ). For RLHF, a separate reward model is usually trained to provide the reward, and then the is fine-tuned to maximize the reward. Though effective, such RLHF methods are even more costly as they require collecting a large amount of annotated data to train an accurate reward model that reflects a desired model behavior. Moreover, once the reward model has been trained, it becomes impossible to add a new behavior control during inference time as cannot be changed on the fly.
In this paper, we take the third choice to maximize , which is to optimize the latent input representations (abbreviated as in the following). Such method is also called “Representation Engineering” proposed in [52], in which authors learn a from contrastive pairs of good and bad samples, which they call “Contrast Vectors”, and then add to each to employ inference-time control. However, such a learning method does not guarantee the learned “Contrast Vectors” can activate the target behavior. We thus propose to learn such “Contrast Vector” () by directly optimizing Eq. 2 using the EM algorithm. Specifically, we treat output as a discrete latent variable that is intractable, which we estimate during the E-step that maximizes suffix score; then during the M-step, we use this variable to optimize H. Specifically:
@@
In E-Step: Use the -th iteration to sample multiple outputs , …, , in which each , then select the best that has highest suffix score.
In M-step: calculate , then , where denotes the step size.
As is shown in Figure 2, using such EM update, at each M-step update we can get a better input query that is able to sample better output that satisfies the behavior constraint (Step 3, highlighted in orange block). We then use such updated query representation to sample new output in E-step (Step 4) that better satisfy the behaviour constraint. Through multi-iteration updates, we can get a final , which is represented as the Key-Value Cache at each Transformer layer, that can be applied to sample all controlled outputs. We call this algorithm SelfControl, with detailed implementation illustrated in Algorithm 1. In practice, we also adjust the step size through line search (search_step_size in Algorithm 2) to ensure the suffix score keeps increasing through multi-iteration.
Compress Suffix Gradients into SelfControl.
SelfControl can efficiently search for proper input representations to enable LLM behavior control at the instance level. However, it is still expensive and requires a gradient-based search for each input during inference time. We thus propose to just use SelfControl to collect a bunch of {(input, )} pairs, and train a Prefix Controller to immediately enable the LM to generate the updated representation. We call this method SelfControl. This idea is motivated by earlier prefix-prompt tuning research [41, 31, 47] that searches for continuous / discrete tokens to maximize certain outputs. Ours is different in that the reward signal as well as the learning objective are collected through the model itself.
The Prefix Controller does not directly change the model parameters and ensures flexible model control. Each Prefix Controller can be treated as a plug-and-play “prompt”, which elicits a certain LM behavior standalone. In addition, by applying multiple such controllers together, we can steer LM output to follow a composite of desired attributes. For example, in Figure 3, we show the model can be guided to behave with more reasoning capability, calmer and more helpful. As is shown in the Figure 3, the Prefix Controller contains two components: a LoRA adapter and a learnable prefix prompt , where and denote the learnable parameters of the adapter and prefix prompt respectively. The learnable prompt is prepended to each input query, while the LoRA adapter serves as an additional module that directly influences the latent representations. Altogether they will affect the latent representations.
Then given the LoRA adapter, prefix prompt, , and input query, we first define as the KV-cache representations with these inputs:
Then, the objective of SelfControl is to train the adapter and prompt to match the generated representation to the representation obtained from regular SelfControl (the , which is highlighted in orange on the right-hand side of Figure 3). In particular, let be a collection of the input-representation pairs, the objective of SelfControl is formulated as the following:
Data filtering.
Specifically, when collecting using SelfControl, we filter the data with two constraints for higher quality: 1) The suffix score of the last output should be larger than the suffix score of the original output by a margin , i.e., , where represents KV-cache from the uncontrolled model; 2) To ensure that the target hidden states are not too drifted away from the original hidden states and affect model fluency, the L2 norm of the target hidden states should be bounded by the scaled L2 norm of the original hidden states, i.e., , where is the proportion constant.
@@ 3 Experiments
In this section, we benchmark SelfControl and SelfControl on various attributes, including emotions, reducing toxicity, helpfulness and harmlessness (HH) dialogue, and reasoning. We elaborate on our evaluation setups below. Further details of our experiments are in Appendix C.

### v3

@@ Abstract
We propose SelfControl, an inference-time model control method utilizing gradients to control the behavior of large language models (LLMs) without explicit human annotations. Given a desired behavior expressed in a natural language suffix string concatenated to the input prompt, SelfControl computes gradients of the LLM’s self-evaluation of the suffix with respect to its latent representations. The gradients are used to directly control the auto-regressive generation process towards desired behaviors, which eliminates human supervision, achieves precise and transparent control, and offers on-the-fly adaptability. To further enhance efficiency, we introduce SelfControl, a compact module that encapsulates the learned representations from gradients into a PrefixController, facilitating efficient inference-time control with no latency compared to the original model and allowing control for multiple behaviors simultaneously. Our experiments demonstrate SelfControl’s efficacy across multiple domains, where it improves over SOTA for 8.3% in detoxification, 3.1% in truthfulness enhancement, 4%10% in controlling on emotion tones, and 48.2% in privacy protection, i.e., completely remove privacy leakage issue. Additionally, we demonstrate that SelfControl can be used for data synthesis and to improve reasoning abilities. Our interactive demo and code are available at Google Colab demo and code.
@@ 1 Introduction
Recent advancements in Large Language Models (LLMs) like GPT (Achiam et al., 2023), Claude (Anthropic, 2023), Gemini (DeepMind, 2023), and LLaMA (Touvron et al., 2023) have showcased capabilities in understanding and generating human-like text. Users can interact with LLMs by writing their instructions in natural language, served as prompt input to LLMs. However, LLMs may often behave in ways that deviate from human intentions (Shen et al., 2023). Such deviations include the generation of toxic content, leakage of private information (Wang et al., 2023), or inappropriate emotional tones (Dathathri et al., 2020). To address these issues, controlling LLMs to ensure alignment with desired behaviors—such as non-toxicity, truthfulness (Ouyang et al., 2022; Dathathri et al., 2020), privacy preservation (Wang et al., 2023), and emotional tone adjustments (Li et al., 2023)—is crucial (see Figure 1).
One approach to enhance LLM control involves fine-tuning model parameters via online (Ouyang et al., 2022) and offline (Rafailov et al., 2024) Reinforcement Learning (RL) methods. However, these methods face several challenges: 1) They demand substantial effort to collect large-scale human-annotated preference data (Bai et al., 2022a); 2) The control process is indirect and lacks transparency, which may fail when encountering out-of-domain behaviors not represented in the preference data (Huang et al., 2023); 3) Incorporating new desired behaviors necessitates additional fine-tuning, which can potentially compromise the control quality of previously aligned behaviors (Qi et al., 2023).
Besides aligning LLM behavior during training time, recently much research has been focused on controlling LLM at inference time, e.g., through latent representation engineering (RepE)(Zou et al., 2023a). These works still require curating a set of contrastive pairs as a demonstration. Building upon this, in this paper, we’re studying whether we can control LLMs’ behavior via their self-evaluation, i.e. use LLMs as a judge to assess the quality of their own outputs (Kadavath et al., 2022; Bai et al., 2022b; Zheng et al., 2023; Yuan et al., 2024). We thus introduce SelfControl, a novel gradient-based framework for precise LLM behavior control.
The core idea of SelfControl is to leverage the LLM’s self-evaluation to control its behavior. For each input prompt, we formulate a desired behavior as a natural language question, asking the model to assess whether its output aligns with the specified behavior. We term the question as suffix string and concatenate the input prompt, the model’s output, and the suffix string, feeding this combined input back into the model. We then compute the likelihood of the behavior-aligned response to the suffix string as a learning signal for behavior control, which we named as suffix score. Then, we compute the suffix score’s gradient with respect to the latent representations of the original input, which term as suffix gradients. These suffix gradients are then utilized to update the latent representations, controlling the LLM’s behavior towards the desired outcome. We run this procedure multiple times to iteratively update the input’s latent representations, each time using the modified representations to generate new model outputs that increasingly align with the desired behavior.
SelfControl offers several advantages over traditional fine-tuning approaches: 1) Elimination of human-annotation: SelfControl leverages the model’s self-evaluation as learning signals, substantially reducing the effort and resources required for preference data collection and scaling. 2) Precise and transparent control: SelfControl operates at inference time and directly modifies the latent representations, which allows for explicit behavior specification and fine-grained control, consequently enhancing control interpretability. 3) On-the-fly adaptability: SelfControl does not alter model parameters, enabling easy implementation of behaviors and control of combinations of multiple behaviors, thus providing unparalleled flexibility. SelfControl demonstrates superior performance compared to contrastive learning-based control methods, particularly in areas such as detoxification, truthfulness enhancement, privacy protection, and emotion control. As illustrated in Figure 1, SelfControl exhibits remarkable flexibility in controlling LLM for multiple attributes simultaneously, e.g., happiness and calmness.
To enhance its adaptability, efficiency and compositionality, we further propose SelfControl on top of SelfControl as a general controller across inputs. The core module of SelfControl is the PrefixController, a prompt-based adapter (Hu et al., 2021; Zhang et al., 2023) optimized to match the latent representations conditioned on this PrefixController to the latent representations under regular SelfControl. PrefixController brings ideas from prefix-prompt tuning research (Shin et al., 2020; Li & Liang, 2021; Yang et al., 2023) to achieve efficient control, which has almost no latency compared to the original model greatly outperforms other control baselines. Furthermore, we show that PrefixController is a learnable and composable module that can be easily integrated into the LLM to control multiple model behaviors simultaneously (e.g., being happier, while staying calm), shown in Figure 1, thereby enhancing the practicality of SelfControl for real-world applications.
In summary, our primary contributions are as follows:
We introduce SelfControl, a gradient-based LLM control framework that leverages the model’s self-evaluation to eliminate the need for human-annotated data, offering more efficient, precise, transparent, and adaptable control.
We further develop SelfControl using PrefixController, a PEFT (parameter-efficient fine-tuning) module that enhances SelfControl’s adaptability and compositionality, enabling the dynamic application of controlling multiple behaviors simultaneously.
We show SelfControl is effective on a diverse range of control tasks to align LLM behaviors with user intentions and ethical standards, including improvements over SOTA by 8.3% in detoxification, 3.1% for truthfulness enhancement, 4%10% for control on emotion tones and 48.2% for privacy protection, i.e., completely remove privacy leakage issue.
2 Related work
LLM Control and Representation Engineering.
Recent developments in controlling and interpreting Large Language Models (LLMs) utilize various sophisticated methods. For behavior control, techniques such as Reinforcement Learning from Human Feedback (RLHF) (Ouyang et al., 2022), Direct Policy Optimization (DPO) (Rafailov et al., 2024), and knowledge editing methods like ROME (Meng et al., 2022a), MEND (Mitchell et al., 2021), and MEMIT (Meng et al., 2022b) modify model outputs or parameters to align with human preferences or factual accuracy. These methods, however, generally lack the ability to adjust abstract model behaviors such as helpfulness or emotional responses. Other strategies, such as Constrained Decoding (CD) (Dathathri et al., 2020), involve constrained optimization and sampling with Langevin dynamics for token-level output control (Kumar et al., 2021; 2022), which may lack flexibility in that they often require human supervision.
Representation Engineering (RepE) (Zou et al., 2023a; Turner et al., 2023; Rimsky et al., 2023) instead, is a flexible method which mainly focus on finding steering vectors to add on LLMs’ hidden representations. It originates from the previous methods that learn to find a direction, e.g. linear probes, and then add/subtract the direction from model hidden representations. Unlike the supervised methods, recent technique such as Activation Addition (Turner et al., 2023) or Contrast Vector (Zou et al., 2023a), directly engineer the steering vector in a zero-shot manner. We start from the point of view that gradients are naturally engineering model representations. Thus, we also treat the gradients constructed in our methods a different kind of steering vectors.
LLM Self Evaluation
LLM self-evaluation has been shown to be effective in answering multi-choice questions (Ren et al., 2023), judging LLMs’ output, and serving as safeguards (Phute et al., 2023). However, some argue that there are some pitfalls (Panickssery et al., 2024; Zheng et al., 2023) in LLM self-evaluation. These pitfalls include position bias, distribution bias (Panickssery et al., 2024), and sycophancy during evaluation. These issues may affect LLMs’ evaluation and lead to undesired consequences. However, they generally do not apply to our method. Similar to Phute et al. (2023), we simply probe LLMs’ next token probability on Yes and No, guiding LLMs toward their own preferences, which has been demonstrated to be feasible by recent study (Yuan et al., 2024).111Due to page limit, please refer to Appendix A for full related works.
@@ 3 SelfControl
In this section, we present our SelfControl framework, which leverages the LLM’s self-evaluation to control its behavior. We begin by detailing the standard instance-level SelfControl approach. This encompasses the process of transforming desired behaviors into suffix strings, computing suffix scores and suffix gradients, and controlling model behaviors through iterative updates to latent representations. Subsequently, we introduce the across-instance version, SelfControl. SelfControl compresses instance-level suffix gradients into a PrefixController, enabling adaptable model control on new inputs and facilitating the simultaneous control of multiple behaviors.
3.1 Instance-Level SelfControl
SelfControl controls the LLM’s by transforming a desired behavior into a natural language question, referred to as a suffix string. The model then performs self-evaluation of its response to this question, generating a suffix score corresponding to the likelihood of the response aligning with the desired behavior. Then, gradients of the suffix score with respect to the latent representations of the original input are computed. Model behaviors are then controlled through iteratively updating the latent representations with the suffix gradients. Figure 2 illustrates this process.
Formally, we consider an -layer autoregressive Transformer-based Language Model with parameters , denoted as . Given a prompt input, such as “The corner store, with your favorite candy, has new ownership.”, we first process it through the model to obtain the latent Key and Value representations for each layer. These representations are denoted as representing the Key and Value matrices for layer , respectively. We use to denote the function for obtaining these latent representations: . Subsequently, we sample an output sequence one token at a time from the model, conditioned on the input representations:
Without any control, the model may generate an undesired output, such as “Oh no! The corner store, where I always get my favorite candy, has new ownership? That’s so sad.”. To perform LLM self-evaluation of the output, we form a suffix string representing the desired behavior. In this case, to improve the output’s happiness, we might use: “Your response was happy. Yes or No?”. Conditioned on this suffix, we probe the probability of the predicted <next-token> being either “Yes” or “No”:
@@ 3.1 Instance-Level SelfControl
Here, “Yes” and “No” are used solely to assess the ’s evaluation of the response for a certain behavior, and are not necessarily the tokens with the highest predicted probability. We quantify the model behavior by defining the suffix score as the probability ratio between “Yes” and “No”:
The suffix score is directly influenced by the output, and output is exactly the object we want to control. A higher score indicates a stronger alignment between the output and the behavior specified in the suffix. Consequently, the objective of LLM control becomes the generation of an output that maximizes this suffix score:
Equation (2) presents three interdependent variables that can potentially be optimized to maximize : output, , and . The most apparent choice is output, which could be directly searched. However, recent research has shown that guided searches of LLM-generated token sequences can be complex and inefficient (Zou et al., 2023b; Huang et al., 2023; Qi et al., 2023; Liu et al., 2023; Wei et al., 2024; Zeng et al., 2024). An alternative approach is to optimize the model parameters , which corresponds to fine-tuning the model, e.g., RLHF. As discussed in the introduction, this approach encounters several challenges, including the need to collect large amounts of human-annotated preference data, a lack of precise control and transparency, and difficulty in incorporating new behaviors during inference time.
Therefore, SelfControl takes the third choice to maximize , which is to optimize the latent input representations (abbreviated as when input is clear from the context). SelfControl computes the suffix gradients and adds to to employ inference-time model control. This optimization process is performed iteratively, with the -th iteration consisting of two steps:
@@
In Step 1: Use the -th iteration to sample multiple outputs , …, , in which each , then select the best that gives the highest .
In Step 2: Calculate , then update , with denoting the step size.
This iterative update process progressively refines the latent representations , enabling the sampling of output in each iteration to increasingly align with the desired behavior. Through multiple iterations, we obtain a final optimized representation , which can be used to sample the controlled outputs at inference time Algorithm 1 provides a detailed pseudocode of SelfControl. In practice, we employ a line search technique to dynamically adjust the step size , ensuring a consistent increase in the suffix score across iterations. (See Algorithm 2 in Appendix F for more details.)
3.2 SelfControl: Compress Suffix Gradients into PrefixController
SelfControl can efficiently search for proper input representations to enable LLM behavior control at the instance level. To further improve it for across-instance control, we propose to compress the suffix gradients from a set of instances runs into a PrefixController, which can be easily integrated into the LLM and used to control the model behavior.
We implement PrefixController as a learnable adapter , which is prepended to each layer of the LLM as a “prefix”, including the input embeddings layer222The soft tokens at the input layer are initialized using a neutral prompt, e.g., ‘‘You are an assistant.’’. Similarly to SelfControl, PrefixController does not directly change the model parameters but control the model through modifying the latent representations at inference time. We denote the latent representations after applying PrefixController as .
To learn the PrefixController, we first run SelfControl to collect a set of pairs. For the best performance, we also filter the dataset by only keeping the instances that have high suffix scores. (See details in Appendix F.). Then we adapt the following objective to minimize the mean squared error between and to optimize the its parameters :
Each learned PrefixController works as an adaptable module that elicits a specific LLM behavior independently. These modules can be used as plug-and-play components to control model behaviors. Furthermore, by combining multiple PrefixController’s, we can guide the LLM output to exhibit a composite of desired behaviors. For instance, as illustrated in Figure 3, we demonstrate that the model can be directed to display increased happiness while maintaining a calmer demeanor.
@@ 4 Experiments
| Attribute | Task Type | Data Source |
| Emotion | Dialogue | Zou et al. (2023a) |
| HH-dialogue | Dialogue | Bai et al. (2022b) |
| Toxicity | Completion | Gehman et al. (2020) |
| Privacy | Completion | Wang et al. (2023) |
| Reasoning | Completion | Cobbe et al. (2021) |
| Truthfulness | ICL | Marks & Tegmark (2023) |
In this section, we evaluate SelfControl and SelfControl on controlling LLM to follow various attributes, including emotions, language detoxification, privacy protection, harmless, helpfulness, reasoning, and in-context learning of truthfulness. Table 1 summarizes the datasets we use. Further details of our experiments are in Appendix C.